OpenStudy (anonymous):

How do i do 1/2a-3=7-3/4a

OpenStudy (anonymous):

2a - 12 = 28 -3a 5a = 40 a = 8

OpenStudy (anonymous):

where are you getting 2a-12 from?

OpenStudy (anonymous):

i multiplied the entire equation by 4

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

how do you multiply the whole thing by 4?

OpenStudy (anonymous):

so i don't have to deal with fractions

OpenStudy (anonymous):

you multiply each term by 4 as long as you do it to every single term on both sides, the equation will remain equal

OpenStudy (anonymous):

wait but what are you multiplying by 4? did you subsatute the fractions by 4??